Assume there is one electron charge -e inside a spherical region of the uniform charge density p (a Thompson hydrogen model)if the sphere of the radius is given is the 1* 10-10 find the frequency of the radiation emitted by the atom
1
Expert's answer
2020-12-21T11:00:19-0500

Answer

This is similiar as Bohr model energy is given by

"E=\frac{-13.6eV}{n^2}"

the frequency of the radiation emitted by the atom

"\nu=\frac{E}{h}"

Using above both equation

"\nu=\frac{E}{h}=\frac{13.6\times1.6\times10^{-19}}{6.6\times10^{-34}}\\=3.3\times10^{15}Hz"